Eve Shift Summary 23 Dec 1999

Ashkan_Alavi-Harati (alavi@fnal.gov)
Fri, 24 Dec 1999 00:14:58 -0600

Shift Summary for evening shift, 23 Dec 99

On Shift: Ashkan, Rickardo

Hours of beam available at all intensities: 6.8/8.5
Hours of beam available at nominal intensity: 6.0
Hours of E799 data (special runs not included): 5.0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

A TeV quench and a ramp dump. ktev2-crash and related daq problems.

========================================================================
Runs ended this shift:

Run 15275 - 4.2M events - BEAM_799

Tapes brought to FCC this shift: none
========================================================================
Accesses during shift: None

========================================================================
Overview of shift: E799 datataking. Some problems with daq especially
ktev2 crash, otherwise relatively smooth running when beam was available.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15275 58 15:31 ktevana errors, reset the pipeline, spills 58-60.

73 15:50 Trigger error, DYC busy, KFREND resolves.

97 16:28 TeV injection quench, estimate 45 min.

123 17:00 Beam is back.

17:15 No trigger since beam returned, busy status
says DDD veto 2 busy, KFRENDing doesn't
help. Bob suggests to do a fresh start.

17:18 Try to stop the run. DAQ is hanging, couldn't
stop the run normally, could not save online
histo's. drt_fresh_start from a new ktev3 window.

18:00 Both daq and sdaq indicate ktev2 is dead. Also
ktev1 is hanging. Page Sydney to interrupt his
Xmas shopping. He kindly directs us through the
phone to reboot ktev2 and check on ktev1.
Rebooting ktev2 apparently solves the ktev1
problem as well!

15279 1 19:10 Finally get the new run going with a fresh start
and brand new tapes. AA updated XT constants,
DC resolutions looked horrible for the last run.
Also updated eop reference plots from run 15275.

19:45 Onsplit tape UPK033 mounted on ktev4.

91 21:14 Tev ramp dump, no beam.

145 22:22 Beam is back.

161 22:45 Trigger error, KFREND.

23:55 Onsplit tape UPF063 mounted on ktev2.

========================================================================
Follow up Notes on Previous Shift

- Could have Taken more good data, had the daq had allowed us.
- Watched channel 968 seems OK. Channel 983 has a 5.5 ped shift.

========================================================================
Items to Follow up on Next Shift

=======================================================================